OpenStudy (anonymous):

Graph the inequality on a coordinate plane. -y≤ 3x-5

OpenStudy (jkristia):

do you know how to grapy -y = 3x - 5 ?

OpenStudy (anonymous):

no

OpenStudy (jkristia):

do you know how to graph a line give y = mx + b ? If so, try change -y = 3x-5 into y = .... format (by multiplying by -1 on both sides)

OpenStudy (anonymous):

i did not anything about this can help me

OpenStudy (jkristia):

do you know how to graph, say, y = 2x + 3 ?

OpenStudy (anonymous):

noooooo ;(

OpenStudy (jkristia):

hmm, you should have learned that before you learn about graphing inequalities. Sorry I cant help. But you can maybe take a look at Khan to see how to graph functions and how to graph a line of the format y = mx + b.